What's the answer for square root of x^2-y^2?
Answer by Theo(13342)

i believe the answer depends on the values of x and y.

if you set z = sqrt(x^2 - y^2), you can square both sides of that equation to get:
z^2 = x^2 - y^2
if you add y^2 to both sides of that equation, then you get:
z^2 + y^2 = x^2 which can be shown as:
x^2 = z^2 + y^2

this looks a lot like the pythagorean formula of c^2 = a^2 + b^2, except:

x = c
z = a
y = b

c is the hypotenuse of a right triangle.
a and b are legs of the right triangle.

this can be translated to:

x is the hypotenuse of a right triangle.
z and y are legs of the right triangle.

is that what you're looking for?

if not, then what i the problem you are trying to solve?